# Human Cognition Inspired RAG with Knowledge Graph for Complex Problem Solving

### v1

@@ 3.2 Mind Map
4 Methods
CogGRAG implements the graph-based RAG paradigm, inspired by human cognitive processes. Specifically, it simulates human cognitive strategies by decomposing complex questions into a structured mind map and incorporates self-verification mechanisms during the reasoning process. The framework of CogGRAG consists of three key steps: decomposition, retrieval, and reasoning with self-verification. The overall workflow of the framework is illustrated in Figure 2.
4.1 Decomposition
When confronted with complex question, humans typically decompose them into a series of simpler sub-questions and address them incrementally. Inspired by this cognitive strategy, in this paper, we decompose the complex question $Q$ in a top-down manner, which are logically interconnected and form a structured mind map. The resulting mind map aids in identifying the necessary information required to answer, even if such information is not explicitly mentioned in the original question. For instance, in the question “The football manager who recruited David Beckham managed Manchester United during what timeframe?”, the key entity “Alex Ferguson” (the manager who recruited David Beckham) might be overlooked by traditional matching-based RAG methods, failing to derive the correct answer.
[…]
$\begin{split}&\left\{\right(q^{t+1}_{j},s^{t+1}_{j}),j=1,...,N\}\\ &=\text{Decompose}(q^{t},p_{\theta},prompt_{1}),\end{split}$ (1)
where $N$ is the number of sub-questions, which is adaptively determined by LLMs based on the logic of each sub-question. $prompt_{1}$ describes the question decomposition task in natural language, as shown in Table 6. After decomposition, each sub-question and its state are added as a new node $(q^{t+1}_{j},t+1,s^{t+1}_{j})$ at level $t+1$. Decomposition is terminated when all the states of the child nodes are “End”, constructing a mind map $M$.
Figure 3: The key information extraction with local and global level.
4.2 Retrieval
At the retrieval step, CogGRAG leverages the mind map generated from question decomposition to retrieve relevant information from an external KG. Firstly, to ensure detailed and comprehensive knowledge integration, we extract key information from all questions in the mind map from two distinct granularity levels to support subsequent retrieval processes. As shown in Figure 3, entities and triples are extracted from each individual question at the local level, while inter-question relationships are captured at the global level and represented as subgraphs. This process can be formalized as follows:
$Keys=\text{Extract}(M,p_{\theta},prompt_{2}),$ (2)
where $Keys$ consists of the sets of entities, triples and subgraphs extracted from the mind map $M$, guided by the extraction workflow outlined in $prompt_{2}$, as shown in Table 7 and Table 8.
Next, based on the extracted key information $Keys$, CogGRAG aims to retrieve relevant structured knowledge (triples) from the KG. Specifically, we expand each entity in the entity set of $Keys$ to its neighbors in the KG, forming triples that link the entity to its neighboring entities via relationships. These retrieved tuples are added to a candidate tuple set $\tilde{\mathcal{T}}$. Subsequently, to reduce redundant and noisy information, we employ the triples and the subgraphs in $Keys$ to prune the triples in $\tilde{\mathcal{T}}$ through similarity matching. In practice, only tuples in $\tilde{\mathcal{T}}$ with similarity scores exceeding a predefined threshold $\varepsilon$ are retained, resulting in the final triples set $\mathcal{T}$. This process can be formalized as follows:
$\begin{split}&\mathcal{T}=\\ &\Big{\{}x\in\tilde{\mathcal{T}}\mid\exists y\in Triple\cup Subgraph,\\ &sim(x,y)>\varepsilon\Big{\}},\end{split}$ (3)
where $Triple\in Keys$ represents the set of triples extracted from individual questions, $Subgraph\in Keys$ denotes the set of subgraphs extracted from inter-question relationships, $sim(\cdot)$ is implemented using cosine similarity, with $\varepsilon$ being a hyper-parameter. This pruning ensures that the tuples retrieved from the KG remain both concise and relevant, enhancing the accuracy of downstream reasoning tasks.
4.3 Reasoning with Self-Verification
During the inference phase, CogGRAG utilizes the triples $\mathcal{T}$ retrieved in the previous step to reason and generate answers for the input question. Drawing upon the self-verification theory Frederick (2005), which posits that humans engage in self-verification to validate their reasoning outcomes during the cognitive process, we design a dual-process framework inspired by dual-process theories Vaisey (2009). To this end, we incorporate two LLMs into this step: one dedicated to the reasoning process and the other to the self-verification of the reasoning results. This approach ensures a more robust and reliable inference mechanism by mimicking the human cognitive process of reasoning followed by critical evaluation.
Specifically, given the triples set $\mathcal{T}$ and the mind map $M$, we prompt the LLM to address all sub-questions in the mind map in a bottom-up manner until reaching the root node, which corresponds to the target question $Q$. At each level $t$, the reasoning model $LLM_{res}$ infers and generates an answer $a^{t}$ for the question $q^{t}$, where $a^{t}=LLM_{res}(\mathcal{T},p_{\theta},q^{t},\hat{M},prompt_{3})$ and $\hat{M}=\left\{(q^{T},\hat{a}^{T}),(q^{T-1},\hat{a}^{T-1}),...,(q^{t+1},\hat{a}^{t+1})\right\}$ comprises all the question-answer pairs that have already been successfully reasoned through. Subsequently, the verification model $LLM_{ver}$ reflects on the current reasoning results $a^{t}$, aiming to identify potential errors or inconsistencies in the reasoning process based on the accumulated reasoning path. If the verification model $LLM_{ver}$ detects an error at the current step $t$, it will prompt the reasoning model $LLM_{res}$ to re-think and generate a new answer $\hat{a}^{t}$, else $\hat{a}^{t}=a^{t}$. Notably, during reasoning, we explicitly prompt the LLM to respond with "I don’t know" for questions it cannot answer, rather than generating incorrect responses. This helps mitigate the hallucination problem commonly observed in LLMs during reasoning tasks. Ultimately, CogGRAG infers the answer $A$ to the target question $Q$, where $A=\hat{a}^{0}$. In summary, reasoning with self-verification employs a bottom-up reasoning process based on the mind map within a dual-process framework. All prompt templates used in CogGRAG can be found in the Appendix D.
Table 1: Datasets statistics.
@@ 4.3 Reasoning with Self-Verification
Type Method HotpotQA CWQ WebQSP
RL EM RL EM RL EM
Without external knowledge graph
LLM-only Direct 19.1% 17.3% 31.4% 28.8% 51.4% 47.9%
CoT 23.3% 20.8% 35.1% 32.7% 55.2% 51.6%
With external knowledge graph
LLM+KG Direct+KG 27.5% 23.7% 39.7% 35.1% 52.5% 49.3%
CoT+KG 28.7% 25.4% 42.2% 37.6% 52.8% 48.1%
Graph-based RAG ToG 29.3% 26.4% 54.6% 49.1% 61.7% 57.4%
MindMap 27.9% 25.6% 46.7% 43.7% 56.6% 53.1%
Ours CogGRAG 34.4% 30.7% 56.3% 53.4% 59.8% 56.1%
5 Experiments
5.1 Experimental Settings
Datasets and evaluation metrics. In order to test CogGRAG’s complex problem-solving capabilities on KGQA tasks, we evaluate CogGRAG on three widely used complex KGQA benchmarks: (1) HotpotQA Yang et al. (2018), (2) WebQSP Yih et al. (2016), (3) CWQ Talmor and Berant (2018). Following previous work Li et al. (2023), full Wikidata Vrandečić and Krötzsch (2014) is used as structured knowledge sources for all of these datasets. Considering that Wikidata is commonly used for training LLMs, there is a need for a domain-specific QA dataset that is not exposed during the pretraining process of LLMs in order to better evaluate the performance. Thus, we also test CogGRAG on a recently released domain-specific dataset GRBENCH Jin et al. (2024b). All methods need interact with domain-specific graphs containing rich knowledge to solve the problem in this dataset. The statistics and details of these datasets can734 be found in Table 1. For all datasets, we use two evaluation metrics: (1) Exact match(EM): measures whether the predicted answer or result matches the target answer exactly. (2) Rouge-L(RL): measures the longest common subsequence of words between the responses and the ground truth answers.
Baselines. In our main results, we compare CogGRAG with three types state-of-the-art methods: (1) LLM-only methods without external knowledge, including direct reasoning and CoT Wei et al. (2022) by LLM. (2) LLM+KG methods integrate relevant knowledge retrieved from the KG into the LLM to assist in reasoning, including direct reasoning and CoT by LLM. (3) Graph-based RAG methods allow KGs and LLMs to work in tandem, complementing each other’s capabilities at each step of graph reasoning, including Mindmap Wen et al. (2023), Think-on-graph Sun et al. (2023) and Graph-CoT Jin et al. (2024b).
Experimental Setup. We conduct experiments with three LLM backbones: LLaMA2-13B Touvron et al. (2023), LLaMA3-8B AI@Meta (2024) and Qwen2.5-7B Yang et al. (2024). For all LLMs, we load the checkpoints from huggingface [^1] https://huggingface.co and use the models directly without fine-tuning. We implemented CogGRAG and conducted the experiments with one A800 GPU [^2] Our code is available at: https://anonymous.4open.science/r/RAG-5883. Consistent with the Think-on-graph settings, we set the temperature parameter to 0.4 during exploration and 0 during reasoning. The threshold $\epsilon$ in the retrieval step is set to $0.7$.
Due to the space limitation, we move details on datasets (Sec A), and baselines (Sec B) to appendices.
5.2 Main Results
We perform experiments to verify the effectiveness of our framework CogGRAG, and report the results in Table 2. We use Rouge-L (RL) and Exact match (EM) as metric for all three datasets. The backbone model for all the methods is LLaMA2-13B. From the table, the following observations can be derived: (1) CogGRAG demonstrates the best results in most cases. (2) Compared to methods that incorporate external knowledge, the LLM-only approach demonstrates significantly inferior performance. This performance gap arises from the lack of necessary knowledge in LLMs for reasoning tasks, highlighting the critical role of external knowledge integration in enhancing the reasoning capabilities of LLMs. (3) Graph-based RAG methods demonstrate superior performance compared to LLM+KG approaches. This performance advantage is particularly evident in complex problems, where not only external knowledge integration but also involving “thinking procedure” is essential. Methods ToG and MindMap exhibit enhanced performance by implementing iterative retrieval and reasoning processes on the KGs, thereby generating the most probable inference outcomes through “thinking” on the KGs.
Table 3: Overall results of our CogGRAG with different backbone models on three KBQA datasets. We highlight the best score on each dataset in bold.
Type Method HotpotQA CWQ WebQSP
RL EM RL EM RL EM
LLM-only Qwen2.5-7B 15.3% 15.0% 25.4% 24.1% 46.7% 45.3%
LLaMA3-8B 17.5% 14.9% 30.3% 27.5% 50.4% 45.1%
LLaMA2-13B 19.1% 17.3% 31.4% 28.8% 51.4% 47.9%
LLM+KG Qwen2.5-7B+KG 24.2% 15.6% 33.8% 32.1% 46.7% 45.3%
LLaMA3-8B+KG 25.9% 21.4% 40.6% 35.3% 53.6% 49.1%
LLaMA2-13B+KG 27.5% 23.7% 39.7% 35.1% 52.5% 49.3%
Graph-Based RAG CogGRAG w/ Qwen2.5-7B 28.4% 27.1% 50.5% 45.7% 53.2% 51.6%
CogGRAG w/ LLaMA3-8B 32.1% 27.2% 53.5% 48.4% 57.2% 55.3%
CogGRAG w/ LLaMA2-13B 34.4% 30.7% 56.3% 53.4% 59.8% 56.1%
Method E-commerce Literature Academic Healthcare
RL EM RL EM RL EM RL EM
LLaMA2-13B 7.1% 6.8% 5.4% 5.1% 5.4% 4.7% 4.3% 3.1%
Graph-CoT 26.4% 24.0% 26.7% 23.3% 19.3% 14.8% 28.1% 25.2%
CogGRAG 30.2% 28.7% 32.4% 30.1% 23.6% 21.5% 27.4% 25.6%
5.3 Performance with different backbone models
We evaluate how different backbone models affect its performance on three datasets HotpotQA, CWQ and WebQSP, and report the results in Table 3. We conduct experiments with three LLM backbones LLaMA2-13B, LLaMA3-8B and Qwen2.5-7B. For all LLMs, we load the checkpoints from huggingface and use the models directly without fine-tuning. From the table, we can observe the following key findings: (1) CogGRAG achieves the best results across all backbone models compared to the baseline approaches, demonstrating the robustness and stability of our method. (2) The performance of our method improves consistently as the model scale increases, reflecting enhanced reasoning capabilities. This trend suggests that our approach has significant potential for further exploration with larger-scale models, indicating promising scalability and adaptability.
5.4 Performance on domain-specific KG
Given the risk of data leakage due to Wikidata being used as pretraining corpora for LLMs, we further evaluate the performance of CogGRAG on a recently released domain-specific dataset GRBENCH Jin et al. (2024b) and report the results in Table 4. This dataset requires all questions to interact with a domain-specific KG. We use Rouge-L (RL) and Exact match (EM) as metrics for this dataset. The backbone model for all the methods is LLaMA2-13B Touvron et al. (2023). The table reveals the following observations: (1) CogGRAG continues to outperform in most cases. This demonstrates that our method consistently achieves stable and reliable results even on domain-specific knowledge graphs. (2) Both CogGRAG and Graph-CoT outperform LLaMA2-13B by more than 20%, which can be ascribed to the fact that LLMs are typically not trained on domain-specific data. In contrast, graph-based RAG methods can effectively supplement LLMs with external knowledge, thereby enhancing their reasoning capabilities. This result underscores the effectiveness of the RAG approach in bridging knowledge gaps and improving performance in specialized domains.
Figure 4: Ablation study on the main components of CogGRAG.
Method Correct Missing Hallucination
LLaMA2-13B 19.1% 25.7% 55.2%
ToG 29.3% 20.2% 50.5%
MindMap 27.9% 22.4% 49.7%
CogGRAG 34.4% 40.6% 44.9%
5.5 Ablation Study
5.6 Hallucination and Missing Evaluation
In the reasoning and self-verification phase of CogGRAG, we prompt the LLM to respond with “I don’t know” when encountering questions with insufficient or incomplete relevant information during the reasoning process. This approach is designed to mitigate the hallucination issue commonly observed in LLMs. To evaluate the effectiveness of this strategy, we test the model on the HotpotQA dataset, with results reported in Table 5. We categorize the responses into three types: “Correct” for accurate answers, “Missing” for cases where the model responds with “I don’t know,” and “Hallucination” for incorrect answers. As shown in the table results, our model demonstrates the ability to refrain from answering questions with insufficient information, significantly reducing the occurrence of hallucinations. This capability highlights the effectiveness of our approach in enhancing the reliability and truthfulness of the model’s responses.
6 Conclusion
In this paper, we proposed a graph-based RAG framework, CogGRAG, to enhance the complex reasoning capabilities of LLMs in KGQA tasks. Inspired by human cognitive processes when tackling complex problems, CogGRAG simulates human reasoning by decomposing complex questions into sub-questions and constructing a structured mind map. Furthermore, recognizing that humans often engage in self-verification during reasoning, CogGRAG incorporates a self-verification mechanism during the reasoning phase, leveraging a dual-process to perform reasoning followed by verification, thereby reducing the occurrence of hallucinations. By integrating decomposition, retrieval, and reasoning with self-verification, CogGRAG significantly improves the accuracy of LLMs in handling complex problems. Extensive experiments demonstrate that CogGRAG outperforms state-of-the-art baselines in the majority of cases.
Limitations
Although CogGRAG significantly enhances the reasoning capabilities of LLMs in complex problems on KGQA tasks, several limitations remain. First, while our method avoids iterative retrieval and reasoning, the processes of problem decomposition and self-verification inevitably introduce additional computational overhead. Thus, there is room for future improvements in optimizing reasoning efficiency for practical applications. Second, CogGRAG is currently limited to graph-based external knowledge and cannot be directly applied to document-based knowledge sources, which restricts its generalizability in real-world scenarios. In the future, we plan to explore methods that integrate multiple types of data structures as external knowledge to further enhance the reasoning capabilities of LLMs. Additionally, research on improving retrieval and reasoning efficiency remains a key focus for our ongoing work.
Ethical Considerations
Our research is fundamental in nature and is not directly linked to specific applications. Consequently, the potential for misuse or negative societal impacts depends on how others may apply our methodology. Furthermore, our work does not involve any stakeholders who may benefit or be disadvantaged, nor does it involve vulnerable groups. All datasets used in this study are publicly available and widely utilized in the research community, ensuring no privacy risks and alignment with their intended purpose for scientific inquiry.
Appendix A Datastes
WebQSP is a dataset designed specifically for question answering and information retrieval tasks, aiming to promote research on multi-hop reasoning and web-based question answering techniques. It contains 4,737 natural language questions that are answerable using a subset Freebase KG Bollacker et al. (2008).

### v2

@@ Introduction
$a^{t}=\texttt{LLM}_{\text{res}}(\mathcal{T},q^{t},\hat{\mathcal{M}},\texttt{prompt}_{res}).$ (4)
Self-Verification and Re-thinking.
The candidate answer $a^{t}$ is passed to the verifier LLM along with the current reasoning path. The verifier assesses consistency, factual grounding, and logical coherence. If the answer fails validation, a re-thinked response $\hat{a}^{t}$ is regenerated:
$\hat{a}^{t}=\begin{cases}\texttt{LLM}_{\text{res}}(\mathcal{T},q^{t},\hat{\mathcal{M}},\texttt{prompt}_{rethink})&\text{if }\delta^{t}=\texttt{False},\\ a^{t}&\text{otherwise.}\end{cases}$ (5)
Here, $\delta^{t}$ is a boolean indicator defined as:
$\delta^{t}=\texttt{LLM}_{\text{ver}}(q^{t},a^{t},\hat{\mathcal{M}},\texttt{prompt}_{ver}),$ (6)
which returns True if the verification module accepts $a^{t}$, and False otherwise. In addition to verification, CogGRAG incorporates a selective abstention mechanism. When the reasoning LLM cannot confidently produce a grounded answer based on $\mathcal{T}$, it is explicitly prompted to respond with “I don’t know” rather than hallucinating. The final answer $A$ to the original question $Q$ is obtained by recursively aggregating verified answers $\hat{a}^{t}$ from the leaf nodes up to the root node $q^{0}$, such that $A=\hat{a}^{0}$.
This bottom-up reasoning pipeline, governed by a globally constructed mind map and protected by verification layers, enables CogGRAG to perform accurate and reliable complex reasoning over structured knowledge. All prompt templates used in CogGRAG can be found in the Appendix.
Type Method HotpotQA CWQ WebQSP
RL EM F1 RL EM F1 RL EM F1
LLM-only Direct 19.1% 17.3% 18.7% 31.4% 28.8% 31.7% 51.4% 47.9% 53.5%
CoT 23.3% 20.8% 22.1% 35.1% 32.7% 33.5% 55.2% 51.6% 55.3%
LLM+KG Direct+KG 27.5% 23.7% 27.6% 39.7% 35.1% 38.3% 52.5% 49.3% 52.1%
CoT+KG 28.7% 25.4% 26.9% 42.2% 37.6% 40.8% 52.8% 48.1% 50.5%
Graph-based RAG ToG 29.3% 26.4% 29.6% 49.1% 46.1% 47.7% 54.6% 57.4% 56.1%
MindMap 27.9% 25.6% 27.8% 46.7% 43.7% 44.1% 56.6% 53.1% 58.3%
RoG 30.7% 28.1% 30.4% 55.3% 51.8% 54.7% 65.2% 62.8% 67.2%
GoG 31.5% 30.1% 31.1% 55.7% 52.4% 54.8% 65.5% 59.1% 63.6%
Ours CogGRAG 34.4% 30.7% 35.5% 56.3% 53.4% 55.8% 59.8% 56.1% 58.9%
Experiments
Experimental Settings
Datasets and evaluation metrics. In order to test CogGRAG’s complex problem-solving capabilities on KGQA tasks, we evaluate CogGRAG on three widely used complex KGQA benchmarks: (1) HotpotQA (yang2018hotpotqa), (2) WebQSP (yih2016value), (3) CWQ (talmor2018web). Following previous work (cok), full Wikidata (wikidata) is used as structured knowledge sources for all of these datasets. Considering that Wikidata is commonly used for training LLMs, there is a need for a domain-specific QA dataset that is not exposed during the pretraining process of LLMs in order to better evaluate the performance. Thus, we also test CogGRAG on a recently released domain-specific dataset GRBENCH (graph_cot). All methods need to interact with domain-specific graphs containing rich knowledge to solve the problem in this dataset. The statistics and details of these datasets can be found in Appendix. For all datasets, we use three evaluation metrics: (1) Exact match (EM): measures whether the predicted answer or result matches the target answer exactly. (2) Rouge-L (RL): measures the longest common subsequence of words between the responses and the ground truth answers. (3) F1 Score (F1): computes the harmonic mean of precision and recall between the predicted and gold answers, capturing both completeness and exactness of overlapping tokens.
Baselines. In our main results, we compare CogGRAG with three types state-of-the-art methods: (1) LLM-only methods without external knowledge, including direct reasoning and CoT (prompt_cot) by LLM. (2) LLM+KG methods integrate relevant knowledge retrieved from the KG into the LLM to assist in reasoning, including direct reasoning and CoT by LLM. (3) Graph-based RAG methods allow KGs and LLMs to work in tandem, complementing each other’s capabilities at each step of graph reasoning, including Mindmap (mindmap), Think-on-graph (Think_on_graph), Graph-CoT (graph_cot), RoG (luo2024rog), GoG (xu2024gog). Due to the space limitation, we move details on datasets, baselines and experimental setup to Appendix.
Experimental Setup. We conduct experiments with four LLM backbones: LLaMA2-13B (llama2), LLaMA3-8B (llama3), Qwen2.5-7B (qwen2) and Qwen2.5-32B (hui2024qwen2). For all LLMs, we load the checkpoints from huggingface [^1] https://huggingface.co and use the models directly without fine-tuning. We implemented CogGRAG and conducted the experiments with one A800 GPU. Consistent with the Think-on-graph settings, we set the temperature parameter to 0.4 during exploration and 0 during reasoning. The threshold $\epsilon$ in the retrieval step is set to $0.7$.
Type Method HotpotQA CWQ WebQSP
RL EM F1 RL EM F1 RL EM F1
LLM-only Qwen2.5-7B 15.3% 15.0% 15.8% 25.4% 24.1% 24.5% 46.7% 45.3% 45.5%
LLaMA3-8B 17.5% 14.9% 16.2% 30.3% 27.5% 29.0% 50.4% 45.1% 48.3%
LLaMA2-13B 19.1% 17.3% 18.7% 31.4% 28.8% 31.7% 51.4% 47.9% 53.5%
Qwen2.5-32B 28.7% 27.4% 28.5% 55.1% 50.3% 54.2% 68.4% 60.5% 65.1%
LLM+KG Qwen2.5-7B+KG 24.2% 15.6% 21.4% 33.8% 32.1% 34.9% 46.7% 45.3% 46.1%
LLaMA3-8B+KG 25.9% 21.4% 23.6% 40.6% 35.3% 39.1% 53.6% 49.1% 52.3%
LLaMA2-13B+KG 27.5% 23.7% 27.6% 39.7% 35.1% 38.3% 52.5% 49.3% 52.1%
Qwen2.5-32B+KG 35.6% 32.8% 34.9% 58.3% 54.7% 57.6% 70.2% 65.1% 68.4%
Graph-Based RAG CogGRAG w/ Qwen2.5-7B 28.4% 27.1% 28.2% 50.5% 45.7% 48.9% 53.2% 51.6% 55.0%
CogGRAG w/ LLaMA3-8B 32.1% 27.2% 31.0% 53.5% 48.4% 52.6% 57.2% 55.3% 55.4%
CogGRAG w/ LLaMA2-13B 34.4% 30.7% 35.5% 56.3% 53.4% 55.8% 59.8% 56.1% 58.9%
CogGRAG w/ Qwen2.5-32B 40.5% 37.1% 40.2% 66.5% 62.7% 65.4% 74.1% 68.3% 73.0%
Table 2: Overall results of our CogGRAG with different backbone models on three KBQA datasets.
Main Results
We perform experiments to verify the effectiveness of our framework CogGRAG, and report the results in Table 1. We use Rouge-L (RL), Exact match (EM) and F1 Score (F1) as metric for all three datasets. The backbone model for all the methods is LLaMA2-13B. From the table, the following observations can be derived: (1) CogGRAG achieves the best results in most cases except on WebSQP. Since the dataset is widely used, we attribute the reason to be data leakage. (2) Compared to methods that incorporate external knowledge, the LLM-only approach demonstrates significantly inferior performance. This performance gap arises from the lack of necessary knowledge in LLMs for reasoning tasks, highlighting the critical role of external knowledge integration in enhancing the reasoning capabilities of LLMs. (3) Graph-based RAG methods demonstrate superior performance compared to LLM+KG approaches. This performance advantage is particularly evident in complex problems, where not only external knowledge integration but also involving “thinking procedure” is essential. These methods synergize LLMs with KGs to enhance performance by retrieving and reasoning over the KGs, thereby generating the most probable inference outcomes through “thinking” on the KGs.
Performance with different backbone models
We evaluate how different backbone models affect its performance on three datasets HotpotQA, CWQ and WebQSP, and report the results in Table 2. We conduct experiments with three LLM backbones LLaMA2-13B, LLaMA3-8B, Qwen2.5-7B and Qwen2.5-32B. For all LLMs, we load the checkpoints from huggingface and use the models directly without fine-tuning. From the table, we can observe the following key findings: (1) CogGRAG achieves the best results across all backbone models compared to the baseline approaches, demonstrating the robustness and stability of our method. (2) The performance of our method improves consistently as the model scale increases, reflecting enhanced reasoning capabilities. This trend suggests that our approach has significant potential for further exploration with larger-scale models, indicating promising scalability and adaptability.
Method E-commerce Literature Academic Healthcare
RL EM F1 RL EM F1 RL EM F1 RL EM F1
LLaMA2-13B 7.1% 6.8% 6.9% 5.4% 5.1% 5.3% 5.4% 4.7% 5.1% 4.3% 3.1% 3.6%
Graph-CoT 26.4% 24.0% 25.3% 26.7% 23.3% 24.9% 19.3% 14.8% 16.9% 28.1% 25.2% 26.7%
CogGRAG 30.2% 28.7% 29.5% 32.4% 30.1% 31.3% 23.6% 21.5% 22.7% 27.4% 25.6% 26.2%
Performance on domain-specific KG
Given the risk of data leakage due to Wikidata being used as pretraining corpora for LLMs, we further evaluate the performance of CogGRAG on a recently released domain-specific dataset GRBENCH (graph_cot) and report the results in Table 3. This dataset requires all questions to interact with a domain-specific KG. We use Rouge-L (RL), Exact match (EM) and F1 Score (F1) as metrics for this dataset. The backbone model for all the methods is LLaMA2-13B (llama2). The table reveals the following observations: (1) CogGRAG continues to outperform in most cases. This demonstrates that our method consistently achieves stable and reliable results even on domain-specific KGs. (2) Both CogGRAG and Graph-CoT outperform LLaMA2-13B by more than 20%, which can be ascribed to the fact that LLMs are typically not trained on domain-specific data. In contrast, graph-based RAG methods can effectively supplement LLMs with external knowledge, thereby enhancing their reasoning capabilities. This result underscores the effectiveness of the RAG approach in bridging knowledge gaps and improving performance in specialized domains.
Ablation Study
Figure 3: Ablation study on the main components of CogGRAG.
Method Correct ($\uparrow$) Missing ($\uparrow$) Hallucination ($\downarrow$)
LLaMA2-13B 19.1% 25.7% 55.2%
ToG 29.3% 20.2% 50.5%
MindMap 27.9% 22.4% 49.7%
CogGRAG 34.4% 40.6% 25.0%
Hallucination and Missing Evaluation
In the reasoning and self-verification phase of CogGRAG, we prompt the LLM to respond with “I don’t know” when encountering questions with insufficient or incomplete relevant information during the reasoning process. This approach is designed to mitigate the hallucination issue commonly observed in LLMs. To evaluate the effectiveness of this strategy, we test the model on the HotpotQA dataset, with results reported in Table 4. We categorize the responses into three types: “Correct” for accurate answers, “Missing” for cases where the model responds with “I don’t know,” and “Hallucination” for incorrect answers. As shown in the table results, our model demonstrates the ability to refrain from answering questions with insufficient information, significantly reducing the occurrence of hallucinations.
Method Academic Amazon Goodreads Disease
Graph-CoT 22.39 29.04 15.79 32.97
GoG 21.50 32.20 18.80 37.50
CogGRAG 18.24 31.57 16.32 35.64
Table 5: Average inference time per question (s).
Performance in Inference Times
We conduct the verification on four datasets GRBENCH-Academic, GRBENCH-Amazon, GRBENCH-Goodreads, and GRBENCH-Disease. Table 5 presents the average runtime of CogGRAG, Graph-CoT and GoG. From the table, we can see that CogGRAG is on par with Graph-CoT and GoG. Although our reasoning process introduces self-verification, which increases the time cost, CogGRAG does not require iterative repeated reasoning and retrieval. Instead, it retrieves all relevant information in one step based on the mind map, avoiding redundant retrieval, which provides an advantage in large-scale KGs. Moreover, while self-verification adds additional time costs, it also improves the accuracy of reasoning results, as confirmed in the ablation experiments.
Conclusion
In this paper, we proposed CogGRAG, a graph-based RAG framework to enhance LLMs’ complex reasoning for KGQA tasks. CogGRAG generates tree-structured mind maps, explicitly encoding semantic relationships among subproblems, which guide both local and global knowledge retrieval. A self-verification module further detects and revises potential errors, forming a dual-phase reasoning paradigm that reduces hallucinations and improves factual consistency. Extensive experiments show that CogGRAG substantially outperforms existing RAG baselines on complex multi-hop QA benchmarks.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 42375146 and National Natural Science Foundation of China No. 62202172.
Appendix A Datastes
WebQSP is a dataset designed specifically for question answering and information retrieval tasks, aiming to promote research on multi-hop reasoning and web-based question answering techniques. It contains 4,737 natural language questions that are answerable using a subset Freebase KG (bollacker2008freebase).
@@ Appendix A Datastes
Table 6: Datasets statistics.
Figure 4: Case of CogGRAG.